Three 1.5 V cells are arranged in series, powering a circuit that has two operations with equivalent resistances of 229 Ω (connecting to the positive terminal) and 393 Ω that meet in a single, metered connection. There is a 17 μC capacitor that bridges between the connection and the negative terminal.

(a) If the potential on the negative side of the battery is -0.5 V, then what is the potential on the positive side of the battery? V (

b) What is the potential at the connection point? V

(c) What is the charge on the capacitor? C
Answer #1

given

V = 1.5 volts

R1 = 229 Ω

R2 = 393 Ω

C = 17 μC

a )

VA = voltage

= - 0.5 V

VB - VA = 1.5 x 3

= 4.5

VB = 4.5 - 0.5

= 4 volts

b )

potential at the connection point A = - 0.5 + 1.5

= 1 volts

potential at the connection point B = 1 + 1.5

= 2.5 volts

c )

the equivalent capacitance Ceq = C1 C2 / C1 + C2

= 17 / 2

= 8.5 μC

Q = C V

= 8.5 x 10-6 x 4.5

Q = 38.25 x 10-6 C

Q = 38.25 μC this is the charge on capacitor in series